Photochronograph Corporation (PC) manufactures time series photographic equipment. It is currently at its target debt-equity ratio of .60. It’s considering building a new $65 million manufacturing facility. This new plant is expected to generate aftertax cash flows of $9.4 million in perpetuity. The company raises all equity from outside financing. There are three financing options:

1.

A new issue of common stock: The flotation costs of the new common stock would be 8 percent of the amount raised. The required return on the company’s new equity is 14 percent.

2.

A new issue of 20-year bonds: The flotation costs of the new bonds would be 4 percent of the proceeds. If the company issues these new bonds at an annual coupon rate of 8 percent, they will sell at par.

3.

Increased use of accounts payable financing: Because this financing is part of the company’s ongoing daily business, it has no flotation costs, and the company assigns it a cost that is the same as the overall firm WACC. Management has a target ratio of accounts payable to long-term debt of .15. (Assume there is no difference between the pretax and aftertax accounts payable cost.)

What is the NPV of the new plant? Assume that PC has a 21 percent tax rate.
Answer #1

WACC = [re x we] + [rd x (1−T)x wd], where

  • re is the cost of equity,
  • rd is the cost of debt,
  • T is the tax rate, and
  • wi is the corresponding market weight, for example we = we / [we + wd]

Since the target Debt / Equity ratio is 0.6, Debt represents 0.6 / 1.6 = 37.5% of the company and Equity represents the remaining 62.5%.

Using the WACC formula,

WACC = [0.14 x 0.625] + [0.08 x (1 - 0.21) x 0.375] = 0.0875 + 0.0237 = 0.1112, or 11.12%

Let,

  • I = initial investment
  • CF = subsequent cash flows
  • r = cost of capital

The net present value (NPV) formula for perpetual cash flows is:

NPV = −I + [CF/r] = −65 + [9.4 / 0.1112] = - 65 + 84.53 = 19.53

The NPV is $19.53 mil